Solve the problem with complete solutions. A brass of 1 kg mass has an initial temperature of 98 degree Celsius is submerged in 0.64 kg of water at 7 degree Celsius. The final temperature is 30 degree Celsius. Find the specific heat capacity of the brass.


1
Expert's answer
2021-12-30T11:54:33-0500
"c_bm_b(t_f-t_{ib})+c_wm_w(t_f-t_{iw})=0,\\\space\\
c_b=\dfrac {c_wm_w(t_{iw}-t_f)}{m(t_f-t_i)}=909\text{ J/(kg°C)}."